OpenStudy (anonymous):

Totally lost...and I hate factorization. Help Please? x^5-5x^4-3x^3 What is the greatest common factor and then have to factor the equation

OpenStudy (tkhunny):

x^3 is a pretty obvious place to start. \(x^{3}(x^{2} - 5x - 3)\) Now what? Note: You don't have an equation. It's an expresion. An "equation" will have one of these "=".

OpenStudy (anonymous):

Its not obvous to me, sorry. i am not sure what exactly to do. Math is not my forte'

geerky42 (geerky42):

Yeah, just like tkhunny said, you can factor x³ out. Now you have x³(x²−5x−3) Can you factor (x²−5x−3)?

geerky42 (geerky42):

If you look closely, you will realize that it cannot be factored, so you are done.

OpenStudy (tkhunny):

Terms are separated by + or -. The idea is to find common elements in each term x^5 = x*x*x*x*x x^4 = x*x*x*x x^3 = x*x*x So, x^3 is common to all

geerky42 (geerky42):

Exactly. Is this clear?

OpenStudy (tkhunny):

Oh, and ... "I hate factorization" "Math is not my forte'" Quit talking like that. It is self-fulfilling. Let it go.

OpenStudy (anonymous):

I see how you got the greatest common factor of x^3 yes, but now i have to factor the original expression ... and i'm sorry tkhunny. it's how i feel. I have never been one to grasp math and i am here to recieve help and advice. Thank you for the help.

geerky42 (geerky42):

"but now i have to factor the original expression" It cannot be factored.

OpenStudy (anonymous):

@geerky42 the question asked me : x^5-5x^4-3x What is the GCF? which you both helped me determine was x^3. The question continues to ask me to factor out the GCF of x^5-5x^4-3x^3=

OpenStudy (tkhunny):

Well, so far, and I have encountered you a time or two, I see no reason for you to feel like that. The most valuable help and advice I could give is to talk you out of this position. The common factors are: x x^2 x^3 Which of these is "the greatest"?

OpenStudy (anonymous):

x^3

OpenStudy (anonymous):

then i'm supposed to factor out the GCF x^5-5x^4=3x^3=

OpenStudy (anonymous):

whooops, x^5-5x^4-3x^3=

OpenStudy (tkhunny):

This was done already. x^5 = x^3 * x^2 -5x^4 = x^3 * (-5x) -3x^3 = x^3 * (-3) Thus x^3 * (x^2 - 5x - 3)

OpenStudy (anonymous):

Thank you for helping me understand. I wrote it down so that it will help me next time.

OpenStudy (tkhunny):

Good call. Give some thought to that other stuff, too. Really, aren't you struggling with enough stuff? Why add math to it? You can significantly reduce your struggles simply by spending no more time, ever, believing that you can't do it.

OpenStudy (anonymous):

Thank you. That's good advice. :)